OpenStudy (anonymous):

While Steve Bouchard was a student at the University of Florida, he borrowed $12,000 in student loans at an annual interest rnte of 9 percent. If Steve repays $1,500 per year, how long to the nearest year will it take him to repay the loan?

OpenStudy (amistre64):

if he pays it before or after the interest rate kicks in, it makes a difference :) maybe. But i gotta quit trying to work thru these on the keyboard ... to many typos ....

OpenStudy (amistre64):

B{n} = 12000(1.09^n) - 1500(1.09^(n+1)-1.09)/.09 B{12} =abt $821.91 left to pay, so at most id say 13 years if he pays 1500 before the interest kicks in.

OpenStudy (amistre64):

If he makes the 1500 payment after interest kicks in we got: B{n} = 12000(1.09^n)-1500(1.09^n -1)/.09 This has a Balance of abt. $1071.95 after 14years sp maybe 15 years to pay off?